# The Discovery of Little Red Dots in the Local Universe: Signatures of Cool Gas Envelopes

### v1

@@ 6.1.4 The Origin: a Cool Gas Envelope around BHs
The unusually high-EW Na D, K I, Fe II, and Ca T absorption lines of The Egg all point to a non-stellar and non-ISM origin. If the optical continuum is AGN-dominated, these absorption lines likely arise from an atmosphere-like gas envelope in the vicinity of the central BHs. Although a stellar population struggles to explain the lines observed in The Egg, the microphysics of line formation may apply to both stellar atmospheres and LRDs. Therefore, a comparison to stellar spectra enables us to infer the physical conditions of the absorbing gas.
@@
In the following, we investigate the EW of Ca T absorption as a key tracer. Figure 12 places The Egg’s CaT absorption in the context of stellar atmospheres. As shown in the left panel of Figure 12, within the stellar catalog presented here, only two metal-rich stars exhibit Ca T EWs greater than that of The Egg. These two stars are classified as an Ia yellow supergiant and a 0–Ia yellow hypergiant, with effective temperature () of 5500 and 5727 K, surface gravities () of 0.0 and 0.4, and metallicities [Fe/H] of and , respectively.
@@
The yellow supergiant and hypergiant stars exhibit the strongest Ca T absorption due to their low surface gravity and their effective temperatures of K. This temperature allows Ca in the atmosphere to exist mainly in the singly-ionized state. In stellar atmospheres, the EWs of Ca T absorption strongly anti-correlate with surface gravity (e.g., S. V. Mallik, 1997; A. J. Cenarro et al., 2002). This is due to the gas opacity changing with the photosphere density, which generally increases with in stellar atmosphere models. As decreases, the photosphere density decreases, leading to a diminished continuum absorption opacity (the main source being H- ions at and at wavelengths around – Å), while the line opacity of Ca II changes only mildly (U. G. Jorgensen et al., 1992). This enhances the contrast between the line and continuum opacities, increasing the EWs of the Ca T absorption. Thus, the Ca T EWs are essentially atmospheric density indicators at a given effective temperature.
@@
We interpret the Ca T absorption in The Egg as originating from an atmosphere-like gas envelope surrounding the BH, with a structure and opacity analogous to that of cool stellar atmospheres. We apply stellar Ca T diagnostics to estimate to first order. The right panel of Figure 12 demonstrates the stellar Ca T EWs as a function of . The Ca T EW– relation also depends on stellar metallicity. Extrapolating the empirical Ca T EW- relation for metal-rich stars with [Fe/H]0.25 to match the observed Ca T EW of The Egg yields an estimated of approximately . If we assume a slightly lower metallicity ( to ), the inferred would be about . The value may be even lower given the low metallicity we have inferred from the NLR (Section 5.2).
@@
In conclusion, the Ca T absorption in The Egg indicates the presence of a metal-rich cool gas envelope with low surface gravity (). Comparison with known supergiant and hypergiant stars suggests that the of such an atmosphere is likely around 5000–6000 K. However, we emphasize that this is a first-order approximation with simplified assumptions; the actual physical conditions of the absorbing gas may differ substantially from those in stellar atmospheres. In particular, the mapping between and the stellar atmosphere density (and hence opacity) is based on hydrostatic equilibrium, which is not necessarily true for the envelope of The Egg. Nevertheless, as we argue above, the line formation process mainly depends on the atmospheric temperature and density. The inferred value of above may not reflect the dynamical condition of the envelope, but does suggest that its atmosphere density may resemble that of very low-surface gravity stars. Standard stellar photosphere relations imply that at K corresponds to a photosphere density of (e.g., C. J. Hansen et al., 2004, their Equations (4.56)(4.65)). Thus, we expect the photosphere density of The Egg to be on the order of or lower than .
@@
We note that the FWHMs of these metal absorbers are about 300–450 km s-1 (Table 2), significantly broader than those typically observed in stellar atmospheres. For instance, Ca T absorption lines in stars are intrinsically narrow, with FWHMs reaching only up to 70 km s-1 in supergiant stars (A. V. Filippenko & L. C. Ho, 2003). The broad Ca T absorption in The Egg suggests that the cool gas envelope is turbulent or dynamically complex. A more detailed analysis, incorporating tailored modeling of the gas properties, is necessary for a comprehensive understanding.
@@
We have limited our analysis to the Ca T lines; similar atmosphere conditions are presumably responsible for the other absorption lines identified here. However, we do not perform the same analysis on them due to our unawareness of dedicated stellar datasets on these lines. Future characterization of observed stellar spectral lines combined with analysis of theoretical spectral libraries can directly test our scenario.
@@ 6.2 Optical to IR Continuum: Evidence of Thermalized Emission
As discussed in §6.1, we find evidence for cool gas envelopes that may be analogous to (supergiant or hypergiant) stellar atmospheres at 5000 K. In this section, we investigate the potential application of atmosphere-like gas envelopes to explain the observed continuum shape of the local LRDs. A similar comparison conducted by A. de Graaff et al. (2025) also noted a resemblance to supergiant stars, though with notable discrepancies. In Figure 13, we compare the spectrum of The Egg to a 5000 K blackbody. The Å regime of The Egg’s spectrum can be reasonably well described by the Rayleigh-Jeans tail of a 5000 K blackbody. However, the Å region shows a significant flux deficit relative to the blue side of the blackbody spectrum, even without subtracting the extrapolation from the UV continuum as a separate component. Varying the blackbody temperature does not yield a better match, as the optical-to-near-IR continuum of The Egg exhibits stronger curvature than a standard blackbody. Such a pseudo-blackbody continuum shape signifies thermalized emission, but may involve intricate physical mechanisms.
@@
Recent theoretical models propose that envelope-like structures can form around SMBHs under super-Eddington accretion and emit thermal radiation with characteristic temperatures of K (D. Kido et al., 2025; H. Liu et al., 2025). To investigate such scenarios, we compare the observed continuum of The Egg with the models from H. Liu et al. (2025), which assume an atmosphere formed by quasi-spherical super-Eddington accretion flows. As shown in Figure 13, one such model from H. Liu et al. (2025) provides a good match to the optical continuum shape of The Egg. This model has a total luminosity and a density of at the photosphere 777 in their notation.; a star at would have this photosphere density at (C. J. Hansen et al., 2004). The photosphere (defined by the Rosseland-mean optical depth) is located at 510 AU with K. The frequency-dependent gas optical depth effects suppress the UV-to-optical flux, making the predicted continuum deviate from a blackbody 888We invoke no dust reddening to the model curve; we have checked that adding a modest level of dust extinction ( mag using the SMC dust extinction law) would further improve the agreement.. Essentially, the optical spectrum is “reddened” by the atmospheric gas; otherwise, we would have to invoke a very steep dust extinction law to explain the optical redness while retaining the declining IR slope (see Appendix D). However, we caution that the model assumes super-Eddington accretion, while current observations do not provide direct evidence of it. The integrated corresponds to about 22% of the Eddington luminosity, based on the listed in Table 1. The true Eddington ratio is unconstrained, as both and the intrinsic bolometric luminosity are uncertain. In addition to the theoretical model, we present an alternative empirical parameterization of the pseudo-blackbody continuum shape in Appendix D.
@@
The origin of the blue UV continuum remains uncertain. Several observations suggest a stellar origin from young host galaxies, based on the lack of UV variability, the absence of high-ionization UV lines, and the extended, complex UV morphologies in a number of high-redshift LRDs (P. Rinaldi et al., 2024; H. B. Akins et al., 2025; W. L. Tee et al., 2025; R. P. Naidu et al., 2025). However, these features do not definitively rule out an AGN origin, either directly from the accretion disk or scattered by electrons/dust (E. Lambrides et al., 2024; Z. Li et al., 2025; K. Inayoshi et al., 2025). This is especially the case when considering the recent detection of AGN-powered high-ionization UV lines in a subset of LRDs (M. Tang et al., 2025). Recently, P. Rinaldi et al. (in prep.) found that a V-shaped continuum similar to that of LRDs originates from the nuclei of a cosmic-noon IR-bright galaxy, further suggesting that AGN emission could contribute to the blue UV continua. At present, we lack spectral information at rest-frame wavelengths below 2000 Å to map diagnostic UV lines that could distinguish between AGN and star-forming activity. As such, the origin of this component remains inconclusive given the current data. Deep UV spectroscopy or polarization measurements, both of which could trace AGN light, is needed to clarify its nature.
@@ 6.3 The Dust Emission
The WISE photometry of local LRDs suggests that the dust content in these objects, while not as strong as implied by a standard AGN dust torus, is also not negligible. A typical AGN dust torus, when scaled to LRD luminosities, would show rapidly increasing emission beyond 1–2 µm, peaking around 10 µm, which is clearly not the case for LRDs (M. Stalevski et al., 2016; B. Wang et al., 2025b). All three local LRDs show flattening or an increase in flux beyond rest-frame 10 µm (W3 and W4 bands; Figure 1). At , this wavelength range is beyond JWST coverage, and it is either only partially covered by Spitzer and Herschel or falls into gaps between their coverage, leaving it largely unconstrained for high-redshift LRDs (H. B. Akins et al., 2024; C. M. Casey et al., 2024, 2025; D. J. Setton et al., 2025).
Qualitatively, the IR excess from the WISE photometry in these local LRDs (Figure 1) reveals at least dust components with temperatures approximately between and K. The emission at rest-frame 3 µm corresponds to the peak of a K blackbody, and the emission at 10 µm matches a K blackbody. As shown in Figure 13, if assuming templates from J. Lyu & G. H. Rieke (2021) that include dust components at , 1038, 300, and 90 K (see Appendix §D.2), the IR emission of The Egg can be well reproduced by significant dust components at 1038 K and 300 K. However, the derived dust properties depend sensitively on the chosen templates. We discuss alternative dust templates based on Haro 11 (J. Lyu et al., 2016) and the degeneracies in Appendix D.3. In the absence of longer-wavelength data, it is difficult to robustly constrain the dust properties, particularly the cold dust content with temperatures below 100 K.
@@ 7 A conceptual picture of LRDs
BH accretion | Broad emission lines (e.g., Balmer and Paschen series, infrared O I, infrared Fe II) with FWHM 1000 km s-1 | Optically thick gas envelope prevents direct observations of BH accretion process. |
AGN Emission Line Regions | Narrow and broad line regions consistent with AGN ionization; [Fe II] emission lines with FWHM and inferred density between broad and narrow components | Modeling of [Fe II] is complex and potentially degenerate |
Cool gas envelope | Metal absorption lines, especially with extremely high-EW Na D, K I, Fe II, and Ca T absorption features in The Egg; the thermalized emission in good agreement with theoretical atmosphere models (H. Liu et al., 2025) | Requires detailed modeling of opacity, cooling, and circum-BH environments; must also account for the FWHMs of absorption lines |
@@
Outflowing optically-thick gas | Blueshifted Balmer and He I absorption lines | Their correlation with other properties of LRDs (i.e., Balmer break strength) remains uncertain; the origin of He I emission (ionized gas versus dense neutral gas) requires further study. |
Dust emission | near-IR excess in WISE photometry (rest-frame 3-20 µm) | Constraints on dust components depend on model assumptions; cannot fully characterize dust properties from the AGN to the ISM due to the lack of longer-wavelength data. |
@@
BLRs and NLRs. The broad Balmer and Paschen lines, along with the infrared O I emission lines (in The Egg, see Figure 2), indicate that the observed line broadening arises from gas motions within the BLRs surrounding the BHs. Analysis of the Balmer decrements, BPT diagrams, and metallicity suggests that the narrow emission lines originate from AGN-powered NLRs with low metallicity and minimal dust.
[Fe II] emitting gas intermediate between the BLRs and NLRs. The AGN-excited [Fe II] emission lines indicate gas with a density of approximately cm-3, intermediate between the BLRs and NLRs. The FWHM of the [Fe II] lines ( 200 km s-1 for The Egg and 300 km s-1 for J1022+0841 in MODS spectra) is likewise between that of the broad ( km s-1) and narrow lines (17 km s-1 for The Egg and 77 km s-1 for J1022+0841 in the Binospec spectra). The absence of these [Fe II] lines in J1047+0739 may be attributed to a lower covering fraction of the intermediate emitting regions.
A cool gas envelope with 5000–6000 K and photosphere density of (). The optical–IR continua of the three local LRDs exhibit pseudo-blackbody shapes indicative of thermalized emission and are in good agreement with theoretical atmosphere models (H. Liu et al., 2025). The presence of extremely high-EW metal absorbers (Na D, K I, Fe II, Ca T) in The Egg further indicates a cool, optically thick gas envelope surrounding the BH. Analysis of the Ca T EW, by analogy with stellar atmospheres, suggests an atmospheric temperature of approximately 5000–6000 K and surface gravity of at the photosphere, corresponding to a density of . The size of this envelope is on the order of AU as inferred from the optical-IR luminosity and the effective temperature. A consequence of an optically thick cool gas envelope is that the BH accretion process, i.e., the emission from the accretion disk itself, is not directly observable.
@@
The viewing angle in this non-spherical geometry can explain the diversity in continuum shapes and the presence or absence of metal absorption features among different objects. Changes in viewing angle can modulate the relative contributions of thermalized emission and AGN photoionized radiation. It also determines whether the line of sight intersects the metal-absorbing layers within the cool gas envelope. For example, the line of sight toward The Egg may intersect the optically thick part of the envelope with high metal-line opacity. This orientation results in a significant contribution to the continuum from the thermalized emission, making the metal absorbers observable. J1022+0841 could be oriented closer to the optically thin region, where the optical depth of metal lines is lower. J1047+0739 could lie even closer to the optically thin region, leading to a reduced contribution from the thermalized emission to the total optical continuum. However, we emphasize that this is a speculative scenario, and a larger statistical sample is needed to better understand the roles of geometry and viewing angle. Differences in the metallicities of the cool gas envelop would also contribute to the diversity of observed metal absorption features.
@@
Optically thick outflows. In most high-redshift LRDs, the Balmer absorption lines are blueshifted, with only a few cases showing redshifted H absorption with minimal velocity offsets (X. Lin et al., 2024; X. Ji et al., 2025; F. D’Eugenio et al., 2025). Up to now, all detected He I absorbers in LRDs are blueshifted (I. Juodžbalis et al., 2024; B. Wang et al., 2025a; F. Loiacono et al., 2025). The Balmer absorbers are superimposed on broad Balmer emission lines, and the He I absorbers are superimposed on both the broad emission and the underlying continuum. We thus attribute these absorbers to optically thick outflowing gas located either within or outside the BLRs. The rare instances of redshifted Balmer absorption with small velocity shifts may indicate gas cooling and inflowing back toward the nucleus. Although the relationship between the absorption and LRD properties (e.g., Balmer break strength) remains unclear, it is a plausible connection that merits further exploration.
@@
Dust at – K. The WISE photometry of the three local LRDs is consistent with dust components with temperatures ranging from to K. These dust components, cooler than typical AGN-heated hot dust but warmer than ISM dust, is likely located outside the 5000 K atmosphere yet still within the AGN vicinity. However, we note that the analysis of the dust components is highly uncertain, model-dependent, and subject to degeneracies.
@@
In the proposed scenario, the BH resides in an environment distinct from that of typical type-1 AGNs. As discussed in §4, it remains unclear whether the empirical BH mass estimators calibrated for type-1 AGNs are valid for LRDs, given their potentially different physical conditions. Assuming a BH mass of –, the corresponding Eddington luminosity is – erg s-1. For the three objects, if we take the total integrated luminosity , assuming that the UV continuum is AGN light, the Eddington ratio would be . The current observations do not provide sufficient evidence to determine whether LRDs are in a super-Eddington accretion mode, as proposed by some theoretical scenarios (e.g., K. Inayoshi et al., 2024; F. Pacucci & R. Narayan, 2024). The accretion rate of the central BH and its disk can significantly influence the geometry of both the disk and the surrounding gas envelope (e.g., C. Dotan & N. J. Shaviv, 2011). We leave a detailed exploration of this to future work.

### v2

@@ 6.1.3 The Origin: a Cool Gas Envelope around BHs
Motivated by the unusually high-EW Na D, K I, and Ca T absorption, along with the simultaneous presence of other cool-star-like low-ionization absorptions in The Egg, we propose that these absorption lines originate from an atmosphere-like gas envelope surrounding the central BH. The microphysics of line formation from stellar atmospheres may likewise apply to this atmosphere-like gas envelope. A comparison with stellar spectra, therefore, enables us to infer the physical conditions of the absorbing gas.
@@
In the following, we investigate the EW of Ca T absorption as a key tracer. Figure 13 places The Egg’s CaT absorption in the context of stellar atmospheres. As shown in the left panel of Figure 13, within the stellar catalog presented here (Mallik11997; Cenarro2001), only two metal-rich stars exhibit Ca T EWs greater than that of The Egg. These two stars are classified as a G0 Ia yellow supergiant and a G4 0–Ia yellow hypergiant, with effective temperature () of 5500 and 5727 K, surface gravities () of 0.0 and 0.4, and metallicities [Fe/H] of and , respectively.
@@
The yellow supergiant and hypergiant stars exhibit the strongest Ca T absorption due to their low surface gravity and their effective temperatures of K. This temperature allows Ca in the atmosphere to exist mainly in the singly-ionized state. In stellar atmospheres, the EWs of Ca T absorption strongly anti-correlate with surface gravity (e.g., Mallik11997; Cenarro2002). This is due to the gas opacity changing with the photosphere density, which generally increases with in stellar atmosphere models. As decreases, the photosphere density decreases, leading to a diminished continuum absorption opacity (the main source being H- ions at and at wavelengths around – Å), while the line opacity of Ca II changes only mildly (Jorgensen1992). This enhances the contrast between the line and continuum opacities, increasing the EWs of the Ca T absorption. Thus, the Ca T EWs are essentially atmospheric density indicators at a given effective temperature.
@@
We interpret the Ca T absorption in The Egg as originating from an atmosphere-like gas envelope surrounding the BH, with a structure and opacity analogous to that of cool stellar atmospheres. We apply stellar Ca T diagnostics to estimate to first order. The right panel of Figure 13 demonstrates the stellar Ca T EWs as a function of . The Ca T EW– relation also depends on stellar metallicity. Extrapolating the empirical Ca T EW- relation for metal-rich stars with [Fe/H]0.25 to match the observed Ca T EW of The Egg yields an estimated of approximately . If we assume a slightly lower metallicity ( to ), the inferred would be about . The value may be even lower given the low metallicity we have inferred from the NLR (§5.2).
@@
In conclusion, the Ca T absorption in The Egg indicates the presence of a metal-rich cool gas envelope with low surface gravity (). Comparison with known supergiant and hypergiant stars suggests that the of such an atmosphere is likely around 5000–6000 K. However, we emphasize that this is a first-order approximation with simplified assumptions; the actual physical conditions of the absorbing gas may differ substantially from those in stellar atmospheres. In particular, the mapping between and the stellar atmosphere density (and hence opacity) is based on hydrostatic equilibrium, which is not necessarily true for the envelope of The Egg. Nevertheless, as we argue above, the line formation process mainly depends on the atmospheric temperature and density. The inferred value of above may not reflect the dynamical condition of the envelope, but does suggest that its atmosphere density may resemble that of very low-surface gravity stars. Standard stellar photosphere relations imply that at K corresponds to a photosphere density of (e.g., Hansen2004, their Equations (4.56)(4.65)). Thus, we expect the photosphere density of The Egg to be on the order of or lower than .
@@
We note that these metal absorbers are offset by to km s-1 from the NLRs, with FWHMs of approximately 300–440 km s-1 (Table 4). It is unclear whether the offset arises from a velocity shift of the NLRs relative to the systemic redshift of the AGNs, or from the expansion of the gas envelope. Velocity offsets of tens to several hundred km s-1 between NLRs and AGN systemic redshifts are common in type-1 AGNs, driven by various dynamical and radiative processes (Bae2014; Shen2016). Meanwhile, the cool gas envelope itself is turbulent and dynamically complex. The FWHMs are significantly broader than those typically observed in stellar atmospheres. For instance, Ca T absorption lines in stars are intrinsically narrow, with FWHMs reaching only up to 70 km s-1 in supergiant stars (Filippenko2003). A more detailed modeling of the gas properties, dynamical state, and the stability of the envelope under the BH accretion activity is necessary for a comprehensive understanding.
@@
We have limited our analysis to the Ca T lines; similar atmospheric conditions are presumably responsible for the other absorption lines identified here. However, we do not perform the same analysis on them due to our unawareness of dedicated stellar datasets on these lines. Future characterization of observed stellar spectral lines, combined with analysis of theoretical spectral libraries, can directly test our scenario.
@@ 6.2 Optical to IR Continuum: Evidence of Thermalized Emission
As discussed in §6.1, we find evidence for cool gas envelopes that may be analogous to (supergiant or hypergiant) stellar atmospheres at 5000 K. In this section, we investigate the potential application of atmosphere-like gas envelopes to explain the observed continuum shape of the local LRDs. A similar comparison conducted by deGraaff2025 also noted a resemblance to supergiant stars, though with notable discrepancies. In Figure 14, we compare the spectrum of The Egg to a 5000 K blackbody. The Å–1.5 µm regime of The Egg’s spectrum can be reasonably well described by the Rayleigh-Jeans tail of a 5000 K blackbody. However, the Å region shows a significant flux deficit relative to the blue side of the blackbody spectrum, even without subtracting the extrapolation from the UV continuum as a separate component. Varying the blackbody temperature does not yield a better match, as the optical-to-near-IR continuum of The Egg exhibits stronger curvature than a standard blackbody. Such a pseudo-blackbody continuum shape signifies thermalized emission, but may involve intricate physical mechanisms.
@@
Theoretical models have proposed that envelope-like structures can form around BHs under super-Eddington accretion and emit thermal radiation (Meier1982a; Meier1982b; Zhou2019; Kido2025; Liu2025). To investigate such scenarios, we compare the observed continuum of The Egg with the models from Liu2025, which assume an atmosphere formed by quasi-spherical super-Eddington accretion flows. As shown in Figure 14, one such model from Liu2025 provides a good match to the optical continuum shape of The Egg. This model has a total luminosity and a density of at the photosphere 888 in their notation.; a star at would have this photosphere density at (Hansen2004). The photosphere (defined by the Rosseland-mean optical depth) is located at 510 AU with K. The frequency-dependent gas optical depth effects suppress the UV-to-optical flux, making the predicted continuum deviate from a blackbody 999We invoke no dust reddening to the model curve; we have checked that adding a modest level of dust extinction ( mag using the SMC dust extinction law) would further improve the agreement.. Essentially, the optical spectrum is “reddened” by the atmospheric gas; otherwise, we would have to invoke a very steep dust extinction law to explain the optical redness while retaining the declining IR slope (see Appendix D). However, we caution that the model of Liu2025 assumes super-Eddington accretion, while current observations do not provide direct evidence of it. The integrated corresponds to about 10% of the Eddington luminosity, based on the listed in Table 2. The true Eddington ratio is unconstrained, as both and the intrinsic bolometric luminosity are uncertain. In addition to the theoretical model, we present an alternative empirical parameterization of the pseudo-blackbody continuum shape in Appendix D.
@@
The origin of the blue UV continuum ( Å) remains uncertain. Several observations suggest a stellar origin, rather than an AGN one, from young host galaxies, based on the lack of UV variability, the absence of high-ionization UV lines, and the extended, complex UV morphologies in a number of high-redshift LRDs (Rinaldi2024; Akins2025; Tee2025; Naidu2025; Zhuang2025). However, these features do not definitively rule out an AGN origin, either directly from the accretion disk or scattered by electrons/dust (Lambrides2024; Li2025; Inayoshi2025b). This is especially the case when considering the recent detection of AGN-powered high-ionization UV lines in a subset of LRDs (Tang2025). Recently, Rinaldi2025 found that a V-shaped continuum, similar to those observed in LRDs, originates from the nuclei of a cosmic-noon IR-bright galaxy. This finding further suggests that AGN emission could contribute to the blue UV continuum, although a circumnuclear starburst on unresolved scales could still play a role. At present, we lack spectral information at rest-frame wavelengths below 2000 Å to map diagnostic UV lines that could distinguish between AGN and star-forming activity (e.g., C IV, He II, N V; Feltre2016; Gutkin2016; Mingozzi2022). As such, the origin of this component remains inconclusive given the current data. Deep UV spectroscopy or polarization measurements, both of which could trace AGN light, are needed to clarify its nature.
@@ 6.3 The Dust Emission
The WISE photometry of local LRDs suggests that the dust content in these objects, while not as strong as implied by a standard AGN dust torus, is also not negligible. A typical AGN dust torus, when scaled to LRD luminosities, would show rapidly increasing emission beyond 1–2 µm, peaking around 10 µm, which is clearly not the case for LRDs (Stalevski2016; BWang2025). All three local LRDs show flattening or an increase in flux beyond rest-frame 10 µm (W3 and W4 bands; Figure 1). At , this wavelength range is beyond JWST coverage, and it is either only partially covered by Spitzer and Herschel or falls into gaps between their coverage, leaving it largely unconstrained for high-redshift LRDs (Akins2024; Casey2024; Casey2025; Setton2025).
Qualitatively, the IR excess from the WISE photometry in these local LRDs (Figure 1) reveals at least dust components with temperatures approximately between and K. The emission at rest-frame 3 µm corresponds to the peak of a K blackbody, and the emission at 10 µm matches a K blackbody. As shown in Figure 14, if assuming templates from Lyu2021 that include dust components at , 1038, 300, and 90 K (see Appendix §D.2), the IR emission of The Egg can be well reproduced by significant dust components at 1038 K and 300 K. However, the derived dust properties depend sensitively on the chosen templates. We discuss alternative dust templates based on Haro 11 (Lyu2016) and the degeneracies in Appendix D.3. In the absence of continuous near- to mid-infrared coverage and far-infrared constraints, it is challenging to robustly constrain the dust structures and grain properties in these systems.
Recent IR and sub-mm observations have placed first constraints on the dust content of high-redshift LRDs (Casey2024; Casey2025; Setton2025; Xiao2025; Li2025; Chen2025). The weak IR emission observed in the three local LRDs in this work is consistent with both the detections and upper limits reported for high-redshift LRDs. Deep ALMA non-detections and stacking analyses indicate modest dust masses () and peak temperatures of 100–300 K (Casey2025; Setton2025) for high-redshift LRDs. These inferred temperatures agree with the dust components detected in the three local LRDs at rest-frame wavelengths µm. For dust attenuation, Chen2025 placed stringent upper limits on high-redshift LRDs (), whereas other studies have inferred non-negligible values () from broad-line Balmer decrements or continuum fitting (Ji2025; D'Eugenio2025_z7p04; Taylor2025), underscoring a tension in current estimates. We do not attempt to precisely quantify the attenuation on the continuum or broad lines in this work. Proper modeling of dust attenuation in such atypical environments requires careful treatment of multiple factors, including the geometry and distribution of dust needed to maintain nearly dust-free NLRs, uncertainties in broad-line Balmer decrements, the survivability of dust grains in BLRs, and the validity of energy balance along the line of sight. We also emphasize the need for longer-wavelength follow-up observations of local LRDs, such as with ALMA. These observations can yield more stringent constraints on the dust content than those currently available for their high-redshift counterparts, pending the arrival of next-generation IR facilities (e.g., PRIMA, Moullet2023).
@@ 7 A conceptual picture of LRDs
| BH accretion | Broad emission lines (e.g., Balmer and Paschen series, infrared O I, infrared Fe II) with FWHM 1000 km s-1 | Optically thick gas envelope prevents direct observations of BH accretion process (e.g., emission directly from the accretion disks). |
| AGN Emission Line Regions | Narrow and broad line regions consistent with AGN ionization; [Fe II] and Fe II emission lines with FWHM and inferred density between broad and narrow components | Modeling of [Fe II] and Fe II is complex and potentially degenerate |
| Cool gas envelope | Cool-star like metal absorption features, especially extremely high-EW Na D, K I, and Ca T absorption lines in The Egg; the thermalized emission in good agreement with theoretical atmosphere models (Liu2025) | Envelopes are dynamically complex, requiring models that capture absorption-line FWHMs and velocities, as well as detailed prescriptions of opacity, cooling, and circum-BH environments. |
@@
| Outflowing/inflowing optically-thick gas | Blueshifted Balmer and He I absorption lines | Their correlation with other properties of LRDs (i.e., Balmer break strength, continuum shape) remains uncertain; the origin of He I emission (ionized gas versus dense neutral gas) requires further study. |
| Dust emission | near- to mid-IR excess in WISE photometry (rest-frame 3-20 µm) | Constraints on dust components depend on model assumptions; cannot fully characterize dust properties from the AGN to the ISM due to the lack of longer-wavelength data. |
@@
BLRs and NLRs. The broad Balmer and Paschen lines, along with the infrared O I µm emission lines (in The Egg, see Figure 2), indicate that the observed line broadening arises from gas motions within the BLRs surrounding the BHs. Analysis of the Balmer decrements, BPT diagrams, and metallicity suggests that the narrow emission lines originate from AGN-powered NLRs with low metallicity and minimal dust. The high decrements of broad Balmer lines (H/H) imply either high-density gas with complex radiative transfer effects or atypical dust properties in BLRs.
[Fe II]-emitting gas intermediate between the BLRs and NLRs. The AGN-excited [Fe II] (or Fe II) emission lines indicate gas with a density of approximately cm-3, intermediate between the BLRs and NLRs. The FWHM of the [Fe II] lines ( 100-300 km s-1) is likewise between that of the broad ( km s-1) and narrow lines ( km s-1). Determining whether these regions can emit other lines requires deeper, higher-resolution spectroscopy to accurately decompose broad lines, such as those in the Balmer and Paschen series, or to detect fainter features. The absence of these [Fe II] lines in J1047+0739 may be attributed to a lower covering fraction of the intermediate emitting regions.
A cool gas envelope with 5000–6000 K and photosphere density of (). The optical–IR continua of the three local LRDs exhibit pseudo-blackbody shapes indicative of thermalized emission and are in good agreement with theoretical atmosphere models (Liu2025). The presence of extremely high-EW metal absorbers (Na D, K I, Ca T) in The Egg, along with other potential G-to-K star-like low-ionization absorbers (Figure 12), further indicates a cool, optically thick gas envelope surrounding the BH. Analysis of the Ca T EW, by analogy with stellar atmospheres, suggests an atmospheric temperature of approximately 5000–6000 K and surface gravity of at the photosphere, corresponding to a density of . The size of this envelope is on the order of AU as inferred from the optical-IR luminosity and the effective temperature. A consequence of an optically thick cool gas envelope is that the BH accretion process, i.e., the emission from the accretion disk itself, is not directly observable. The origin of the UV emission remains inconclusive with the current data, and we do not attempt to address it within this framework. It could originate either from stars in the host galaxy or from AGN light scattered by the surrounding gas. We leave a more thorough investigation of its nature to future work and observations.
@@
The viewing angle in this non-spherical geometry can explain the diversity in continuum shapes and the presence or absence of metal absorption features among different objects. Changes in viewing angle can modulate the relative contributions of thermalized emission and AGN photoionized radiation. It also determines whether the line of sight intersects the metal-absorbing layers within the cool gas envelope. For example, the line of sight toward The Egg may intersect the optically thick part of the envelope with high metal-line opacity. This orientation results in a significant contribution to the continuum from the thermalized emission, making the metal absorbers observable. J1022+0841 could be oriented closer to the optically thin region, where the optical depth of metal lines is lower. J1047+0739 could lie even closer to the optically thin region, leading to a reduced contribution from the thermalized emission to the total optical continuum. However, we emphasize that this is a speculative scenario, and a larger statistical sample is needed to better understand the roles of geometry and viewing angle. Differences in the metallicities of the cool gas envelope would also contribute to the diversity of observed metal absorption features.
@@
Optically thick outflows/inflows. In most high-redshift LRDs, the Balmer absorption lines are blueshifted, with only a few cases showing redshifted H absorption with minimal velocity offsets (Lin2024; Ji2025; DEugenio2025). Up to now, all detected He I absorbers in LRDs are blueshifted (Juodzbalis2024; Wang2025; Loiacono2025). The Balmer absorbers are superimposed on broad Balmer emission lines, and the He I absorbers are superimposed on both the broad emission and the underlying continuum. We thus attribute these absorbers to optically thick outflowing gas located either within or outside the BLRs. Balmer absorption arises from high-density neutral gas. He I absorbers may originate from either high-density neutral gas clouds responsible for the Balmer absorption or partially ionized gas clouds with lower density. The rare instances of redshifted Balmer absorption with small velocity shifts, like those seen in The Egg and J1022+0841, may indicate inflows, which could represent accreting material or gas cooling and moving back toward the nucleus. The relationship between these absorptions and other properties of LRDs (e.g., Balmer break strength, continuum shape) is not yet established. However, since their presence can be attributed to high-density hydrogen ( cm-3; Inayoshi2024), there may be plausible connections that merit further investigation.
@@
Dust at – K. The WISE photometry of the three local LRDs is consistent with dust components with temperatures ranging from to K. These dust components, cooler than typical AGN-heated hot dust but warmer than ISM dust, are likely located outside the 5000 K atmosphere yet still within the AGN vicinity. However, we note that the analysis of the dust components is highly uncertain, model-dependent, and subject to degeneracies. Due to the lack of observational constraints, the exact dust geometry and grain properties remain unknown. Future modeling efforts with improved IR observations are needed.
@@
In the proposed scenario, the BH resides in an environment distinct from that of typical type-1 AGNs. As discussed in §4, it remains unclear whether the empirical BH mass estimators calibrated for type-1 AGNs are valid for LRDs, given their potentially different physical conditions. Assuming a BH mass of –, the corresponding Eddington luminosity is – erg s-1. For the three objects, if we take the total integrated luminosity , assuming that the UV continuum is AGN light, the Eddington ratio would be . The current observations do not provide sufficient evidence to determine whether LRDs are in a super-Eddington accretion mode, as proposed by some theoretical scenarios (e.g., Inayoshi2024a; Pacucci2024; Madau2025). The accretion rate of the central BH and its disk can significantly influence the geometry of both the disk and the surrounding gas envelope (e.g., Dotan2011). We leave a detailed exploration of this to future work.